### v1

@@ Abstract
Missing entries in multi-dimensional data pose significant challenges for downstream analysis across diverse real-world applications. These data are naturally modeled as tensors, and recent completion methods integrating global low-rank priors with plug-and-play denoisers have demonstrated strong empirical performance. However, these approaches often rely on empirical convergence alone or unrealistic assumptions, such as deep denoisers acting as proximal operators of implicit regularizers, which generally does not hold. To address these limitations, we propose a novel tensor completion framework grounded in the monotone inclusion paradigm, which unifies generalized low-rank priors with deep pseudo-contractive denoisers and extends beyond traditional convex optimization. Building on the Davis–Yin splitting scheme, we develop the GTCTV-DPC algorithm and rigorously establish its global convergence. Extensive experiments demonstrate that GTCTV-DPC consistently outperforms existing methods in both quantitative metrics and visual quality, particularly at low sampling rates.
keywords:
tensor completion, monotone inclusion, generalized low-rank priors, deep pseudo-contractive denoisers
{MSCcodes}
65K10, 68T07, 94A08
1 Introduction
The presence of missing entries in data, often resulting from sensor malfunctions, occlusions, or transmission errors, poses a persistent challenge in data analysis [29]. Tensors, as multi-dimensional arrays, provide a versatile framework for modeling diverse real-world datasets, such as multi-/hyper-spectral images (MSI/HSI) [44], color videos [36], and spatio-temporal traffic data [6]. Consequently, tensor completion has become a pivotal research area, attracting significant attention within the scientific community [39, 43]. In this work, we address tensor completion from highly incomplete observations by developing a unified framework grounded in the monotone inclusion paradigm, which integrates generalized low-rank priors with deep denoising priors.
@@ 1 Introduction
1.1 Related works
1.1.1 Low-rank based methods
Low-rankness has emerged as a dominant prior for designing $R$, as it facilitates the extraction of essential structures from high-dimensional data [30]. One prominent approach involves tensor decomposition, such as CANDECOMP/PARAFAC (CP) decomposition [23], Tucker decomposition [45], and tensor train decomposition [35], to represent low-rank properties, with regularization applied to the decomposed components [8].
Alternatively, tensor nuclear norm (TNN)-based methods, which serve as surrogates for tensor rank, have been shown to outperform decomposition-based approaches [22, 28]. For multi-dimensional image (MDI) completion, many methods rely on the tubal nuclear norm or its variants derived from tensor singular value decomposition (t-SVD) [24]. For instance, Jiang et al. [22] employ a framelet transform-based t-SVD and propose a related TNN for visual tensor completion. Additionally, Wang et al. [47] extend this concept by applying TNN to tensor gradients, yielding a tensor correlated total variation (t-CTV) regularizer that jointly exploits low-rankness and smoothness, thereby enhancing completion efficiency. From a functional perspective, Wang et al. [46] generalize t-SVD to functional transforms, resulting in a functional TNN that captures both global low-rank structure and local smoothness. In traffic data completion, Tucker rank approximations via Tucker decomposition remain popular [11, 34].
Recently, deep learning has been leveraged to learn latent low-rank representations through neural architectures. For example, Luo et al. [32] propose a low-rank tensor function representation (LRTFR) parameterized by multi-layer perceptrons based on Tucker decomposition, while Su et al. [44] develop a deep fully-connected tensor network decomposition to capture fine details. Li et al. [26] introduce a simplified deep rank-one tensor functional factorization (DRO-TFF), although its self-supervised performance is limited.
1.1.2 Multi-prior methods
@@
For instance, Zhao et al. [56] combine FFDNet [55] with TNN for MDI completion, achieving significant improvements through the learned prior but providing only empirical convergence. Zhao et al. [57] extend this by incorporating both FFDNet and BM3D [14] denoisers within the alternating direction method of multipliers (ADMM) [5] framework and prove convergence under the assumption that $\operatorname{D}_{\sigma}$ is the proximal operator of a Kurdyka–Łojasiewicz (KL) regularizer [2], i.e., $\operatorname{D}_{\sigma}=\operatorname{Prox}_{R}$. However, recent works [21, 51] demonstrate that deep denoisers are generally not conservative and thus cannot be represented as proximal mappings. Moreover, enforcing denoisers to be proximal often imposes restrictive Lipschitz assumptions, such as residual or firm nonexpansiveness [21, 38], which can compromise denoising performance. Additionally, verifying the KL property of the underlying prior $R$ is challenging in practice. Liu et al. [28] pair FFDNet with a weighted TNN for anomaly detection in remote sensing images via the ADMM framework, but the convergence analysis still assumes denoisers are proximal operators of implicit regularizers. Liang et al. [27] combine FFDNet and BM4D [33] denoisers, proving fixed-point convergence for a multi-block PnP-ADMM algorithm under bounded denoiser assumptions. However, this fixed point may not correspond to the solution of any optimization objective, limiting interpretability. For traffic data imputation, Chen et al. [6] incorporate a nonconvex tensor low-rank prior and the deep PnP denoiser DRUNet [54], providing only empirical convergence analysis for the multi-prior completion method.
1.2 Contributions
@@
The synergy between global low-rank priors and deep denoisers has shown considerable potential in tensor completion tasks. However, the limitations highlighted above reveal a gap: low-rank priors effectively capture global structure but often overlook fine local details, whereas optimization-based integrations of deep priors typically assume that $\operatorname{D}_{\sigma}$ acts as a proximal operator of an implicit regularizer. Such an assumption typically requires the denoiser to be conservative and to satisfy restrictive properties (e.g., firmly nonexpansive), conditions that are difficult to verify in practice and often detrimental to denoising performance. This motivates a paragidm shift from classical optimization formulations to a broader monotone inclusion framework, in which priors are treated as general operators rather than as proximal maps. Within this framework, we can couple weakly convex low-rank regularizers with deep denoisers that satisfy more natural operator properties (e.g., pseudo-contractive), while still obtaining rigorous convergence guarantees.
Our key contributions are summarized as follows:
@@
$\bullet$ Tensor completion via monotone inclusion. We propose a novel tensor completion model within the monotone inclusion framework, integrating with generalized tensor correlated total variation (GTCTV) and deep pseudo-contractive (DPC) denoisers. Building on the Davis-Yin splitting (DYS), we derive the GTCTV-DPC algorithm in Algorithm 2.
$\bullet$ Rigorous global convergence analysis. We provide a theoretical convergence guarantee for GTCTV-DPC in Corollary 4.4. Starting with the general monotone inclusion problem, we prove that the associated operator $\operatorname{T}$ in Eq. Eq. 8 is strictly pseudo-contractive and extend the admissible stepsize range for DYS in Theorem 4.1.
@@
$\bullet$ Comprehensive empirical validations. Extensive experiments demonstrate that GTCTV-DPC consistently outperforms existing methods in both quantitative metrics and qualitative visual quality, particularly under low sampling rates.
1.3 Paper organization
@@
The remainder of this paper is arranged as follows. Section 2 introduces notations and preliminaries. Section 3 presents the proposed tensor completion method within the monotone inclusion framework. Section 4 provides the convergence analysis. Section 5 reports numerical experiments validating the proposed method. Finally, Section 6 provides a brief conclusion. The supplement material is available at the Google drive [^1] https://drive.google.com/drive/folders/1MSR91ZgSeEJwFrZEDl1yyqGSjKMGh2ra.
2 Notations and preliminaries
@@
Table 1: Notations used in this paper.
@@
2.1 Notations
In this paper, we adopt the tensor notation from [48, 40] and focus on order-$N$ tensors with $N\geq 3$. The primary tensor space under consideration is denoted by $\mathbb{V}=\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{N}}$, with $\mathcal{O}\in\mathbb{V}$ denoting the zero tensor. Additional notations are summarized in Table 1.
@@ 2.1 Notations
Definition 2.1 (tensor-tensor product [40]).
@@
2.2 Relevant concepts of functional analysis
@@
Definition 2.3 (convex functions [4]).
@@
(1) $f(\theta x+(1-\theta)y)\leq\theta f(x)+(1-\theta)f(y).$
Definition 2.4 ($\mu$-weakly convex functions [15]).
@@
Definition 2.5 (firmly nonexpansive [3]).
@@
Definition 2.7 ($\alpha$-averaged [3]).
An operator $\operatorname{D}:V\to V$ is $\alpha$-averaged, for some $\alpha\in(0,1)$, if $\operatorname{D}=(1-\alpha)\operatorname{Id}+\alpha\operatorname{N}$, where $\operatorname{N}:V\to V$ is nonexpansive.
Definition 2.8 (pseudo-contractive (PC) [3]).
@@
(2) $\|\operatorname{D}x-\operatorname{D}y\|^{2}\leq\|x-y\|^{2}+k\|(\operatorname{Id}-\operatorname{D})x-(\operatorname{Id}-\operatorname{D})y\|^{2}.$
@@
Note that, firm nonexpansiveness is a special case of $\alpha$-averaged with $\alpha=0.5$, and non-expansiveness is a special case of pseudo-contractivity with $k=0$. The relationships among these properties are as follows [50]:
(3) $\text{firmly nonexpansive}\implies\alpha\text{-averaged}\implies\text{nonexpansive}\implies\text{PC}.$
Specifically, Wei et al. [50] established a useful connection between $k$-SPC and nonexpansive operators, summarized in Lemma 2.9. This lemma is crucial for providing an alternative perspective on Theorem 4.1, as further discussed in Remark 4.6.
Lemma 2.9 (Lemma 2.1 in [50]).
$\operatorname{D}:V\to V$ is $k$-SPC with $k\in(0,1)$, if and only if, $\operatorname{D}$ can be expressed as
$\operatorname{D}=\frac{1}{1-k}\operatorname{N}-\frac{k}{1-k}\operatorname{Id},$
where $\operatorname{N}$ is nonexpansive.
Definition 2.10 ($L$-Lipschitz [3]).
An operator $\operatorname{D}:V\to V$ is $L$-Lipschitz for $L>0$, if for any $x,y\in V$,
(4) $\|\operatorname{D}x-\operatorname{D}y\|\leq L\|x-y\|.$
Definition 2.11 ($\beta$-cocoercive [3]).
@@
(5) $\langle\operatorname{D}x-\operatorname{D}y,x-y\rangle\geq\beta\|\operatorname{D}x-\operatorname{D}y\|^{2}.$
@@
Lemma 2.12 (Moreau, Theorem 20.25 in [3]).
Let $f:V\to(-\infty,+\infty]$ be a proper, closed, and convex function. Then $\partial f$ is maximally monotone.
Moreover, according to Example 23.3 in [3], when $f:V\to(-\infty,\infty]$ is proper, closed, and convex, and $\tau>0$, then for any $x\in V$, we have
(6) $\operatorname{J}_{\tau\partial f}(x)=\operatorname{Prox}_{\tau f}(x):=\arg\inf_{z\in V}\left(f(z)+\frac{1}{2\tau}\|z-x\|^{2}\right).$
2.3 Monotone inclusion problem
@@
To the best of our knowledge, recent multi-prior tensor-recovery methods [56, 57, 28, 27, 6] remain rooted in the classical optimization paradigm and typically treat the denoiser $\operatorname{D}_{\sigma}$ as a proximal map, solving the resulting problems via ADMM-type schemes. As discussed in Section 1.1.2, their convergence guarantees are largely empirical or rely on restrictive and often impractical assumptions about the denoiser [21, 51]. Liang et al. [27] establish fixed-point convergence for a multi-block PnP-ADMM under a bounded-denoiser assumption. However, the resulting fixed point generally does not correspond to the solution of any optimization objective, which limits interpretability. Figure 1 illustrates the distinction between existing multi-prior methods and our proposed approach: by formulating the problem within the monotone inclusion framework, we treat priors as general operators rather than proximal mappings, thereby imposing substantially fewer constraints on $\operatorname{D}_{\sigma}$.
Figure 1: Schematic comparison of recent multi-prior (optimization-based) methods and our monotone-inclusion-based approach. The latter treats priors as general operators and therefore requires fewer constraints on the denoiser $\operatorname{D}_{\sigma}$.
In this paper, we focus on the following MIP with three operators:
(7) $\text{find }x\in V,\text{ such that }0\in\operatorname{A}x+\operatorname{B}x+\operatorname{C}x,$
@@
where $\operatorname{A},\operatorname{B},\operatorname{C}$ are maximally monotone operators, and $\operatorname{C}$ is additionally $\beta$-cocoercive. To solve the MIP in Eq. Eq. 7, Davis and Yin [16] proposed the well-known Davis-Yin splitting (DYS) method. They reframe the MIP as a fixed-point problem:
$\text{find }z\in V\text{, such that }z=\operatorname{T}z,$
where $\operatorname{T}$ with stepsize $\tau$ is defined as
(8) $\operatorname{T}:=\operatorname{Id}-\operatorname{J}_{\tau\operatorname{B}}+\operatorname{J}_{\tau\operatorname{A}}\circ(2\operatorname{J}_{\tau\operatorname{B}}-\operatorname{Id}-\tau\operatorname{C}\circ\operatorname{J}_{\tau\operatorname{B}}).$
The concrete algorithm of DYS, which employs a Krasnoselskii-Mann (KM) type iteration [20, 41], is outlined in Algorithm 1. When $\tau\in(0,2\beta)$, Davis and Yin [16] proved that the operator $\operatorname{T}$ in Eq. Eq. 8 is $\kappa$-averaged with $\kappa=\frac{2\beta}{4\beta-\tau}$, ensuring convergence of the KM iterations to a fixed point. More recently, Aragón-Artacho and Torregrosa-Belén [1] extended this result, establishing convergence over the broader interval $\tau\in(0,4\beta)$. In contrast to these prior analyses, we show that $\operatorname{T}$ is SPC when $\tau\in(0,4\beta)$ and, by invoking Lemma 10.6 in [12], we further establish global convergence under different choices of the relaxation parameters $\{\lambda_{t}\}_{t\geq 0}$; details can be found in Theorem 4.1.
Algorithm 1 DYS Algorithm for Solving the MIP in Eq. Eq. 7.
Input: $z_{0}$, $\tau$, $N_{\text{max}}$, $\{\lambda_{t}\}_{t=0}^{N_{\text{max}}}$;
for $t=0:N_{\text{max}}-1$ do
  $z_{t+1}=z_{t}+\lambda_{t}\left(x_{t+1}^{\operatorname{A}}-x_{t+1}^{\operatorname{B}}\right)$. // $z_{t+1}=(1-\lambda_{t})z_{t}+\lambda_{t}\operatorname{T}z_{t}$.
end for
 Return $x_{t+1}^{\operatorname{A}}$.
3 Tensor completion via monotone inclusion

### v2

@@ Abstract
Missing entries in multi-dimensional data pose significant challenges for downstream analysis across diverse real-world applications. These data are naturally represented as tensors, and recent completion methods integrating global low-rank priors with plug-and-play denoisers have demonstrated strong empirical performance. However, these approaches often rely on empirical convergence alone or unrealistic assumptions, such as deep denoisers acting as proximal operators of implicit regularizers, which generally does not hold. To address these limitations, we propose a novel tensor completion framework grounded in the monotone inclusion paradigm. Within this framework, deep denoisers are treated as general operators that require far fewer restrictions than in classical optimization-based formulations. To better capture holistic structure, we further incorporate generalized low-rank priors with weakly convex penalties. Building upon the Davis–Yin splitting scheme, we develop the GTCTV-DPC algorithm and rigorously establish its global convergence. Extensive experiments demonstrate that GTCTV-DPC consistently outperforms existing methods in both quantitative metrics and visual quality, particularly at low sampling rates. For instance, at a sampling rate of 0.05 for multi-dimensional image completion, GTCTV-DPC achieves an average mean peak-signal-to-noise ratio (MPSNR) that surpasses the second-best method by 0.717 dB, and 0.649 dB for multi-spectral images, and color videos, respectively.
I Introduction
The presence of missing entries in data, often resulting from sensor malfunctions, occlusions, or transmission errors, poses a persistent challenge in data analysis [1]. Tensors, as multi-dimensional arrays, provide a versatile framework for modeling diverse real-world datasets, such as color images [2], multi-/hyper-spectral images (MSI/HSI) [3], color videos [4], and spatio-temporal traffic data [5]. Consequently, tensor completion has become a pivotal research area, attracting significant attention within the scientific community [6, 7, 8]. In this work, we address tensor completion from highly incomplete observations by developing a unified framework grounded in the monotone inclusion paradigm, which integrates generalized low-rank priors with deep denoising priors.
@@ I Introduction
Figure 1: Schematic comparison of recent multi-prior (optimization-based) methods and the proposed monotone-inclusion-based approach. Unlike the former, which treat the denoiser $\operatorname{D}\sigma$ as a proximal mapping under restrictive assumptions, our framework regards deep priors as general operators, imposing much weaker constraints on $\operatorname{D}\sigma$. This relaxation enables more effective denoising [9] and ultimately yields superior restoration quality.
I-A Related works
I-A1 Low-rank based methods
Low-rankness has emerged as a dominant prior for designing $R$, as it facilitates the extraction of essential structures from high-dimensional data [10]. One prominent approach involves tensor decomposition, such as CANDECOMP/PARAFAC (CP) decomposition [11], Tucker decomposition [12], and tensor train decomposition [13], to represent low-rank properties, with regularization applied to the decomposed components [14, 15, 16, 17, 18, 19, 20].
Alternatively, tensor nuclear norm (TNN)-based methods, which serve as surrogates for tensor rank, have been shown to outperform decomposition-based approaches [21, 22, 23]. For multi-dimensional image (MDI) completion, many methods rely on the tubal nuclear norm or its variants derived from tensor singular value decomposition (t-SVD) [24]. For instance, Jiang et al. [21] employ a framelet transform-based t-SVD and propose a related TNN for visual tensor completion. Additionally, Wang et al. [25] extend this concept by applying TNN to tensor gradients, yielding a tensor correlated total variation (t-CTV) regularizer that jointly exploits low-rankness and smoothness, thereby enhancing completion efficiency. From a functional perspective, Wang et al. [26] generalize t-SVD to functional transforms, resulting in a functional TNN that captures both global low-rank structure and local smoothness. In traffic data completion, Tucker rank approximations via Tucker decomposition remain popular [27, 28, 29]. For instance, Nie et al. [30] introduce a truncated Schatten $p$-norm (TSpN) to mitigate over-shrinkage of TNN, paired with an adaptive truncation rate decay strategy to handle varying missing rates.
Recently, deep learning, both supervised and unsupervised, has been leveraged to learn latent low-rank representations through neural architectures [31, 32, 33]. However, supervised methods require intricate training and often lack generalization, performing poorly on diverse datasets without retraining [34, 35]. In unsupervised methods, Luo et al. [36] propose a low-rank tensor function representation (LRTFR) parameterized by multi-layer perceptrons based on Tucker decomposition, while Su et al. [3] develop a deep fully-connected tensor network decomposition to capture fine details. Li et al. [37] introduce a simplified deep rank-one tensor functional factorization (DRO-TFF), although its self-supervised performance is limited.
I-A2 Multi-prior methods
@@
For instance, Zhao et al. [38] combine FFDNet [39] with TNN for MDI completion, achieving significant improvements through the learned prior but providing only empirical convergence. Zhao et al. [40] extend this by incorporating both FFDNet and BM3D [41] denoisers within the alternating direction method of multipliers (ADMM) [42] framework and prove convergence under the assumption that $\operatorname{D}_{\sigma}$ is the proximal operator of a Kurdyka–Łojasiewicz (KL) regularizer [43], i.e., $\operatorname{D}_{\sigma}=\operatorname{Prox}_{R}$. However, recent works [44, 45] demonstrate that deep denoisers are generally not conservative and thus cannot be represented as proximal mappings. Moreover, enforcing denoisers to be proximal often imposes restrictive Lipschitz assumptions, such as residual or firm nonexpansiveness [44, 46], which can compromise denoising performance. Additionally, verifying the KL property of the underlying prior $R$ is challenging in practice. Liu et al. [23] pair FFDNet with a weighted TNN for anomaly detection in remote sensing images via the ADMM framework, but the convergence analysis still assumes denoisers are proximal operators of implicit regularizers.
Liang et al. [47] integrate FFDNet and BM4D [48] denoisers with a fiber-rank constraint and establish fixed-point convergence for a multi-block PnP-ADMM algorithm under bounded-denoiser assumptions. Hu et al.[49] integrate two pretrained deep neural networks, one for image denoising and another for completion, with TNN for noisy tensor completion, assuming the networks are nonexpansive and proving fixed-point convergence of their algorithm. Nevertheless, this fixed point may not align with the solution to any optimization objective, thereby limiting interpretability. For traffic data imputation, Chen et al. [50] incorporate a nonconvex tensor low-rank prior and the deep PnP denoiser DRUNet [51], providing only empirical convergence analysis for the multi-prior completion method.
I-B Contributions
@@
The synergy between global low-rank priors and deep denoisers has shown considerable potential in tensor completion tasks. However, the limitations highlighted above reveal a gap: low-rank priors effectively capture global structure but often overlook fine local details, whereas optimization-based integrations of deep priors typically assume that $\operatorname{D}_{\sigma}$ acts as a proximal operator of an implicit regularizer. Such an assumption typically requires the denoiser to be conservative and to satisfy restrictive properties (e.g., firmly nonexpansive), conditions that are difficult to verify in practice and often detrimental to denoising performance.
This motivates a paragidm shift from classical optimization formulations to a broader monotone inclusion framework, in which priors are treated as general operators rather than as proximal maps. Within this framework, we can couple weakly convex low-rank regularizers with deep denoisers that satisfy more natural operator properties (e.g., pseudo-contractive), while still obtaining rigorous convergence guarantees. Fig. 1 illustrates the distinction between existing multi-prior methods and our proposed approach. Our key contributions are summarized as follows:
@@
  • Tensor completion via monotone inclusion. We propose a novel tensor completion model formulated within the monotone inclusion framework. The model integrates a generalized tensor correlated total variation (GTCTV) prior to capture global structural dependencies, while incorporating deep pseudo-contractive (DPC) denoisers to preserve fine local details, thereby enhancing overall reconstruction accuracy.
  • Rigorous global convergence analysis. Building upon the Davis–Yin splitting (DYS) scheme, we develop the GTCTV-DPC algorithm in Algorithm 2 and provide a theoretical proof of its global convergence in Corollary 8.
@@
  • Comprehensive empirical validation. Extensive experiments demonstrate that GTCTV-DPC consistently outperforms existing methods in both quantitative metrics and visual quality, particularly at low sampling rates. For instance, at a sampling rate of 0.05 for MDI completion, GTCTV-DPC achieves an average MPSNR that surpasses the second-best method by 0.717 dB, and 0.649 dB for MSIs, and color videos, respectively.
@@
The remainder of this paper is arranged as follows. Section II introduces notations and preliminaries. Section III presents the proposed tensor completion method within the monotone inclusion framework, and provides a rigorous convergence analysis. Section IV reports numerical experiments validating the proposed method. Finally, section V provides a brief conclusion.
II Notations and preliminaries
@@
TABLE I: Notations used in this paper.
@@
II-A Notations
In this paper, we adopt the tensor notations from [25, 52, 53] and focus on order-$N$ tensors with $N\geq 3$. The primary tensor space under consideration is denoted by $\mathbb{V}=\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{N}}$, with $\mathcal{O}\in\mathbb{V}$ denoting the zero tensor. Additional notations are summarized in Table I.
@@ II-A Notations
Definition 1 (tensor-tensor product [52]).
@@
II-B Relevant concepts of functional analysis
@@
Definition 2 (convex functions [55]).
@@
$f(\theta x+(1-\theta)y)\leq\theta f(x)+(1-\theta)f(y).$ (1)
Definition 3 ($\mu$-weakly convex functions [56]).
@@
Definition 4 (firmly nonexpansive [54]).
@@
Definition 6 (pseudo-contractive (PC) [54]).
@@
$\|\operatorname{D}x-\operatorname{D}y\|^{2}\leq\|x-y\|^{2}+k\|(\operatorname{Id}-\operatorname{D})x-(\operatorname{Id}-\operatorname{D})y\|^{2}.$ (2)
@@
Note that nonexpansiveness is a special case of pseudo-contractivity with $k=0$. The relationships among these operator properties are summarized as follows [9]:
$\text{firmly nonexpansive}\;\implies\;\text{nonexpansive}\;\implies\;\text{PC},$ (3)
indicating that pseudo-contractivity imposes the weakest restriction among them.
Definition 7 ($\beta$-cocoercive [54]).
@@
$\langle\operatorname{D}x-\operatorname{D}y,x-y\rangle\geq\beta\|\operatorname{D}x-\operatorname{D}y\|^{2}.$ (4)
@@
Lemma 2 (Theorem 20.25 and Example 23.3 in [54]).
Let $f:V\to(-\infty,+\infty]$ be a proper, closed, and convex function. Then its subdifferential operator $\partial f$ is maximally monotone. Moreover, for any $\tau>0$, the resolvent of $\tau\partial f$ coincides with the proximal operator of $\tau f$, that is, for any $x\in V$,
$\operatorname{J}_{\tau\partial f}(x)=$ $\operatorname{Prox}_{\tau f}(x)$ (5)
= $\arg\min_{z\in V}\left\{f(z)+\frac{1}{2\tau}\|z-x\|^{2}\right\}.$
II-C Monotone inclusion problem
@@
To the best of our knowledge, recent multi-prior tensor-recovery methods [38, 40, 23, 47, 50, 49] remain rooted in the classical optimization paradigm and typically treat the denoiser $\operatorname{D}_{\sigma}$ as a proximal map, solving the resulting problems via ADMM-type schemes. As discussed in section I-A2, their convergence guarantees are largely empirical or rely on restrictive and often impractical assumptions about the denoiser [44, 45]. Liang et al. [47] established fixed-point convergence for a multi-block PnP-ADMM under a bounded-denoiser assumption, while Hu et al. [49] proved fixed-point convergence for PnP-ADMM under a nonexpansive assumption on deep neural networks. However, the resulting fixed point generally does not correspond to the solution of any optimization objective, which limits interpretability. Fig. 1 illustrates the distinction between existing multi-prior methods and our proposed approach: by formulating the problem within the monotone inclusion framework, we treat priors as general operators rather than proximal mappings, thereby imposing substantially milder constraint on $\operatorname{D}_{\sigma}$. In this paper, we focus on the following MIP with three operators:
$\text{find }x\in V,\text{ such that }0\in\operatorname{A}x+\operatorname{B}x+\operatorname{C}x,$ (6)
@@
where $\operatorname{A},\operatorname{B},\operatorname{C}$ are maximally monotone operators, and $\operatorname{C}$ is additionally $\beta$-cocoercive.
III Method